OpenStudy (anonymous):

Solve The Equation.. sine(2(theta))=cosine(theta)

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ sin(2\theta)\implies 2sin(\theta)cos(\theta)}}\\ \quad \\ \quad \\ {\color{blue}{ sin(2\theta)}}=cos(\theta)\) expand and solve for \(\theta\)